OpenStudy (anonymous):

∑1∞(-1)^n/n^1/3 determines the series whether converges conditionally, converges absolutely or diverges, justify your answer.

OpenStudy (mathmate):

Does the series \[\sum_{n=1}^{\infty}|1/n^{\frac{1}{3}}|\] converge? If it does it converges absolutely, otherwise, it converges conditionally at best. (Hint: this is a p-series and convergence depends on the power of n)

OpenStudy (anonymous):

Use the alternating series test. For a series of the form \(\sum (-1)^na_n\), the series converges if \[(1)~~~~a_n\text{ is a decreasing sequence}\\ (2)~~~~a_n\to0\] Additionally, the series absolutely converges if \(\sum |(-1)^na_n|\) also converges.

OpenStudy (anonymous):

tq